OpenStudy (anonymous):

Solve: 2x + 36 + 40 = 180

OpenStudy (blurbendy):

2x + 36 +40 = 180 2x + 76 = 180 2x = 104 x = 104 / 2 Can you solve from here?

OpenStudy (anonymous):

2x+36+40=180 2x+76=180 2x=104 x=52

OpenStudy (hba):

Giving direct answers is not allowed on OS @jbrarman Please check the CoC- http://openstudy.com/code-of-conduct Thanks in advance Happy Openstudying :)

OpenStudy (anonymous):

Really? Is mine allowed since I explained with the answer?